For the population of 18 74 year old females in the United States height is normally distributed with mean 63.9 inches and standard deviation 2.6. If we randomly select a woman from the population what’s the probability that she’s less than 60 inches


1
Expert's answer
2022-06-13T14:30:04-0400

"P(X<60)=P(Z<\frac{60-63.9}{2.6})=P(Z<-1.5)=1-P(Z<1.5)=0.0668."